# Incoherent Approximation of Leakage in Quantum Error Correction

### v1

@@ Abstract
Quantum error correcting codes typically do not account for quantum state transitions - leakage - out of the computational subspace. Since these errors can last for multiple detection rounds they can significantly contribute to logical errors. It is therefore important to understand how to numerically model them efficiently. Fully quantum simulations of leakage require more levels per leaked qubit, which substantially limits the system sizes that may be simulated. To address this, we introduce a Random Phase Approximation (RPA) on quantum channels that preserves the incoherence between the computational and leakage subspaces. The assumption of incoherence enables the quantum simulation of leakage at little computational overhead. We motivate the approximation’s validity by showing that incoherence is achieved naturally during repeated stabilizer measurements. Additionally, we provide various simulation results which show that the RPA yields accurate error correction statistics in the repetition and surface codes with physical error parameters.
@@ I Introduction
In addition to posing a challenge to error correction, leakage also dramatically increases the cost of simulating such experiments. While there do exist efficient classical simulators of many error correcting codes relying on the stabilizer representation gidney2021stim ; aaronson-gottesman-stabilizer – which can accommodate thousands of qubits – certain noise models (such as leakage) do not fit naturally into this framework. This motivates the need for fully quantum simulations for use in verification and validation. However, a fully quantum simulation including leakage becomes severely limited in the system sizes that can be addressed at reasonable computational cost. For example, the memory requirements for such a simulation of $n$ three-level systems scales as $O(3^{n})$ compared to $O(2^{n})$ in the qubit case. Given a fixed budget, this effectively reduces the achievable system size $n$. As an example, the distance 5 rotated surface code bombin-rotated-SC ; Horsman_2012 involves 25 data qubits, which for qutrits would be equivalent to a 40 qubit simulation. In this paper we discuss a general method for approximating leakage as an incoherent process. The incoherence of transitions between the computational and leaked subspaces means that a given qubit will never be in superposition of more than two basis states. This allows for quantum simulations with leakage (3 or 4 levels per qudit), but only using the resources of a qubit (2 level) system. The approximation can be derived as a Random Phase Approximation (RPA) at the level of noisy quantum channels [Sect. II.4]. (Note, our use of the term RPA is independent from the original one e.g. bohm-rpa .)
While the RPA, in principle, can be applied to any quantum circuit, it is particularly amenable to error correction circuits with repetitive stabilizer measurements. Such measurements cause the coherence between the qubit and leakage subspace to decay exponentially in the number of error correction rounds [App. C]. Such an approximation is exact in the case of incoherent leakage (such as from certain thermal channels). It is less accurate when the origin of leakage is coherent (such as during two-qubit diabatic gates), though we find that the accuracy depends strongly on the circuit layout and physical qubit parameters. For realistic parameters (including both coherent and incoherent leakage for three-level systems) based on recent experiments scaling_SC_Google_2022 , the RPA faithfully reproduces the full qutrit simulation statistics in repetition (up to distance 9) and surface codes (distance 3) [App. G]. Interestingly, even in the case where leakage is strongly coherent and the parameters chosen are such that the naive application of the RPA fails to reproduce the exact simulation, one can still derive an effective incoherent model. The effective model accurately predicts the ideal simulation statistics, such as leakage populations and logical error rates [Sect. III]. This is particularly useful, as we require leakage to be incoherent to efficiently simulate large scale error correction experiments in the stabilizer formalism scaling_SC_Google_2022 .
@@ II.2 Representation of classical and quantum channels
We can also generalize the quantum channels description to include classical registers. Each classical register can be formally represented as an independent subsystem. The defining property of a classical subsystem is that the state vector is always incoherent with respect to a fixed basis $\{|j\rangle\}$. (For a quantum trajectory, this means that the subsystem always remains in a product state.) While incoherence is discussed in general detail in Sect. II.3, for the case of classical registers this means that the state vector is always supported by exactly one basis operator. In other words, at any given time in the trajectories simulation, exactly one observable of the set $\{|j\rangle\!\langle j|\}$ has non-vanishing expectation value. Since the full system will always be of tensor product form
@@
Channels acting on both classical and quantum degrees of freedom can be expressed in terms of Kraus operators of the form
$K_{j}^{(r)}\otimes|r\rangle\!\langle r|,$
@@
$K_{j}\otimes|j\rangle$
where $K_{j}$ acts on the quantum degrees of freedom and $|j\rangle$ denotes the value of a new register. These Kraus operators can be used to model the recording of a sampling outcome (i.e., which index $j$ was sampled) into a classical register. A typical example is the recording of a POVM $\{M_{j}\}_{j}$, which can be represented using Kraus operators $\sqrt{M_{j}}\otimes|j\rangle$.
@@ II.3 Simulating incoherent systems
II.4 Random Phase Approximation
@@
which assigns a distinct phase $\phi_{n}$ to each subspace. The Random Phase Approximation (RPA) is equivalent twirling over the set of unitaries $U(\bar{\phi})$ dankert2009exact . This means conjugating with respect to $U(\bar{\phi})$ and its inverse and averaging over independent and uniformly distributed phases $\phi_{n}$,
$\mathcal{E}\rightarrow\mathcal{E}_{RPA}=\langle\hat{U}(-\bar{\phi})\circ\mathcal{E}\circ\hat{U}(\bar{\phi})\rangle_{\bar{\phi}}.$ (9)
@@
$\mathcal{E}_{RPA}(\rho)$ $=\sum_{j}\langle U(-\bar{\phi})K_{j}U(\bar{\phi})\,\rho\,U^{\dagger}(\bar{\phi})K_{j}^{\dagger}U^{\dagger}(-\bar{\phi})\rangle_{\bar{\phi}}$
@@
where the second term corresponds to II and the first to I and not II. The Kraus operators generated under the RPA can take two possible forms. First there are cross terms
@@
We note some important properties of the RPA. First, although the Kraus operators generated in Eqs. (10) and (11) depend on the specific choice of Kraus operators $K_{j}$ used to represent $\mathcal{E}$, the resulting channel in (9) is representation independent. Thus, if $\{K_{j}\},\{K_{j}^{\prime}\}$ are two sets of distinct Kraus representations of $\mathcal{E}$, the resulting Kraus operators from equations (10), (11) will correspond to the same channel. Second, observe that both forms of Kraus operators for $\mathcal{E}_{RPA}$ satisfy Eq. (5). Therefore $\mathcal{E}_{RPA}$ is an incoherent channel for every $\mathcal{E}$. In fact, $\mathcal{E_{RPA}}$ satisfies an even stronger condition: Inserting the Kraus operators of $\mathcal{E}_{RPA}$ into (5), we see that for each fixed $j$, the function $m(j,n)$ is injective in $n$ (i.e., each output has at most one unique input). This classifies $\mathcal{E}_{RPA}$ as a strictly incoherent operation incoh-ops . This class of operations is a strict subset of the incoherent operations Q-coherence-review . Third, we note that if one of the projectors $P_{n}=P_{C}$ represents the full computational subspace and $\mathcal{E}$ denotes an error channel, then the RPA preserves the process fidelity of that channel
$F_{\text{pro}}(\mathcal{E}_{RPA},\mathcal{I})=F_{\text{pro}}(\mathcal{E},\mathcal{I})\,.$ (12)
@@
where $D=2^{k}$ is the size of the (local) computational subspace on which the channel acts. Finally, we point out that the RPA is a weaker approximation than simply dephasing the system,
@@
since RPA preserves the identity channel.

### v2

@@ Abstract
Quantum error correcting codes typically do not account for quantum state transitions - leakage - out of the computational subspace. Since these errors can last for multiple detection rounds they can significantly contribute to logical errors. It is therefore important to understand how to numerically model them efficiently. Fully quantum simulations of leakage require more levels per leaked qubit, which substantially limits the system sizes that may be simulated. To address this, we introduce a Subspace Twirling Approximation (STA) on quantum channels that preserves the incoherence between the computational and leakage subspaces. The assumption of incoherence enables the quantum simulation of leakage at little computational overhead. We motivate the approximation’s validity by showing that incoherence is achieved naturally during repeated stabilizer measurements. Additionally, we provide various simulation results which show that the STA yields accurate error correction statistics in the repetition and surface codes with physical error parameters.
@@ I Introduction
In addition to posing a challenge to error correction, leakage also dramatically increases the cost of simulating such experiments. While there do exist efficient classical simulators of many error correcting codes relying on the stabilizer representation gidney2021stim ; aaronson-gottesman-stabilizer – which can accommodate thousands of qubits – certain noise models (such as leakage) do not fit naturally into this framework. This motivates the need for fully quantum simulations for use in verification and validation. However, a fully quantum simulation including leakage becomes severely limited in the system sizes that can be addressed at reasonable computational cost. For example, the memory requirements for such a simulation of $n$ three-level systems scales as $O(3^{n})$ compared to $O(2^{n})$ in the qubit case. Given a fixed budget, this effectively reduces the achievable system size $n$. As an example, the distance 5 rotated surface code bombin-rotated-SC ; Horsman_2012 involves 25 data qubits, which for qutrits would be equivalent to a 40 qubit simulation. In this paper we discuss a general method for approximating leakage as an incoherent process. The incoherence of transitions between the computational and leaked subspaces means that a given qubit will never be in superposition of more than two basis states. This allows for quantum simulations with leakage (3 or 4 levels per qudit), but only using the resources of a qubit (2 level) system. The approximation can be derived as a Subspace Twirling Approximation (STA) at the level of noisy quantum channels [Sect. II.4].
While the STA, in principle, can be applied to any quantum circuit, it is particularly amenable to error correction circuits with repetitive stabilizer measurements. Such measurements cause the coherence between the qubit and leakage subspace to decay exponentially in the number of error correction rounds [Sec. III.1]. For realistic models and parameters based on recent experiments scaling_SC_Google_2022 , the STA faithfully reproduces the full qutrit simulation statistics in repetition (up to distance 9) and surface codes (distance 3) [App. F]. In the case where the leakage mechanism is purely coherent, using the STA results in a discrepancy with exact simulation. Even so, we show that one can derive an effective thermal leakage model which is both consistent with coherent leakage and for which the STA is an accurate approximation. The effective model accurately predicts the ideal simulation statistics, such as leakage populations and logical error rates [Sect. III]. This is particularly useful, as we require leakage to be incoherent to efficiently simulate large scale error correction experiments in the stabilizer formalism scaling_SC_Google_2022 .
@@ II.2 Representation of classical and quantum channels
We can also generalize our quantum channels description to include classical registers. This generalization (see the Kraus operators in (2) and (3) below) allows us to model quantum instrumentsdavies1970operational . These are useful in contexts where classical information is required to describe the evolution of quantum systems. Examples include quantum communicationchitambar2014everything , error correctioncharacterizingquantuminstruments , and quantum metrologymclaren2023stochastic ; mitra2022compatibility .
The defining property of a classical register is that, as a local subsystem, it is always incoherent with respect to a fixed basis $\{|j\rangle\}$. (For a quantum trajectory, this means that the subsystem always remains in a product state.) While incoherence is discussed in general detail in Sect. II.3, for the case of classical registers this means that the state vector is always supported by exactly one basis operator. In other words, at any given time in the trajectories simulation, exactly one observable of the set $\{|j\rangle\!\langle j|\}$ has non-vanishing expectation value. Since the full system will always be of tensor product form
@@
Quantum channels conditioned on classical registers can be expressed in terms of Kraus operators of the form
$K_{j}^{(r)}\otimes|r\rangle\!\langle r|,$ (2)
@@
$K_{j}\otimes|j\rangle$ (3)
where $K_{j}$ acts on the quantum degrees of freedom and $|j\rangle$ denotes the value of a new register. These Kraus operators can be used to model the recording of a sampling outcome (i.e., which index $j$ was sampled) into a classical register. A typical example is the recording of a POVM $\{M_{j}\}_{j}$, which can be represented using Kraus operators $\sqrt{M_{j}}\otimes|j\rangle$. Channels of this form are referred to as “fine-grained” quantum instrumentschitambar2014everything .
@@ II.3 Simulating incoherent systems
II.4 Subspace Twirling Approximation
@@
which assigns a distinct phase $\phi_{n}$ to each subspace. The Subspace Twirling Approximation (STA) is equivalent twirling over the set of unitaries $U(\bar{\phi})$ dankert2009exact ; bartlett2007reference . This means conjugating with respect to $U(\bar{\phi})$ and its inverse and averaging over independent and uniformly distributed phases $\phi_{n}$,
$\mathcal{E}\rightarrow\mathcal{E}_{STA}=\langle\hat{U}_{-\bar{\phi}}\circ\mathcal{E}\circ\hat{U}_{\bar{\phi}}\rangle_{\bar{\phi}},$ (11)
where $\hat{U}_{\bar{\phi}}$ is the quantum channel associated with the unitary $U(\bar{\phi})$.
@@
$\mathcal{E}_{STA}(\rho)$ $=\sum_{j}\langle U(-\bar{\phi})K_{j}U(\bar{\phi})\,\rho\,U^{\dagger}(\bar{\phi})K_{j}^{\dagger}U^{\dagger}(-\bar{\phi})\rangle_{\bar{\phi}}$
@@
where the second term corresponds to II and the first to I and not II. The Kraus operators generated under the STA can take two possible forms. First there are cross terms
@@
We note some important properties of the STA. First, although the Kraus operators generated in Eqs. (12) and (13) depend on the specific choice of Kraus operators $K_{j}$ used to represent $\mathcal{E}$, the resulting channel in (11) is representation independent. Thus, if $\{K_{j}\},\{K_{j}^{\prime}\}$ are two sets of distinct Kraus representations of $\mathcal{E}$, the resulting Kraus operators from equations (12), (13) will correspond to the same channel. Second, observe that both forms of Kraus operators for $\mathcal{E}_{STA}$ satisfy Eq. (7). Therefore $\mathcal{E}_{STA}$ is an incoherent channel for every $\mathcal{E}$. In fact, $\mathcal{E_{STA}}$ satisfies an even stronger condition: Inserting the Kraus operators of $\mathcal{E}_{STA}$ into (7), we see that for each fixed $j$, the function $m(j,n)$ is injective in $n$ (i.e., each output has at most one unique input). This classifies $\mathcal{E}_{STA}$ within the strictly incoherent operations incoh-ops , which is a strict subset of the incoherent operations Q-coherence-review . Third, we note that if one of the projectors $P_{n}=P_{C}$ represents the full computational subspace and $\mathcal{E}$ denotes an error channel, then the STA preserves the process fidelity of that channel
$F_{\text{pro}}(\mathcal{E}_{STA},\mathcal{I})=F_{\text{pro}}(\mathcal{E},\mathcal{I})\,.$ (14)
@@
where $D=2^{k}$ is the size of the (local) computational subspace on which the channel acts. Additionally, we point out that the STA is a weaker approximation than simply dephasing the system,
@@
since STA preserves the identity channel. Finally, we comment on the connection of the STA to twirlingbartlett2007reference . The unitary channel $\hat{U}_{\bar{\phi}}$ is a representation of $U(1)^{\times m}$, where $m$ is the number of distinct subspaces $\mathcal{H}_{n}$. The twirling operation represented by equation (11) ensures that $\mathcal{E}_{STA}$ is a morphism of this representation, which means that it commutes with $\hat{U}_{\bar{\phi}}$ for every phase $\bar{\phi}$. This guarantees that $\mathcal{E}_{STA}$ preserves incoherence of states: observe that a state is incoherent with respect to the subspaces $\mathcal{H}_{n}$ if and only if it is invariant under $\hat{U}_{\bar{\phi}}$, i.e.
$\hat{U}_{\bar{\phi}}(\rho)=\rho$
for all $\bar{\phi}$. Because $\mathcal{E}_{STA}$ commutes with $\hat{U}_{\bar{\phi}}$, we have that
$\mathcal{E}_{STA}(\rho)=(\mathcal{E}_{STA}\circ\hat{U}_{\bar{\phi}})(\rho)=(\hat{U}_{\bar{\phi}}\circ\mathcal{E}_{STA})\,(\rho),$
and is therefore still an incoherent state.